# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 8 Appendix - Proofs
By Assumptions A1A4, the covariates are bounded, $\|{\bf Z}\|\leq B_{{\bf Z}}$ a.s., and the hazards are uniformly bounded. Then, there exist $0<c_{h}<C_{h}<\infty$ such that
$c_{h}\;\leq\;\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\;\leq\;C_{h}$
for all $\eta$ in the sieve and all $(t,{\bf X},{\bf Z})$. Hence the integrand in $J(\eta;{\bf W})$ is uniformly bounded by
$\|{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\|\;\leq\;B_{{\bf Z}}^{2}C_{h}$
for all $\eta\in\mathcal{N}_{n}$ and all $\bf W$}. Thus the class $\mathcal{J}_{n}=\bigl\{J(\eta;\cdot):\eta\in\mathcal{N}_{n}\bigr\}$ has a finite envelope and inherits the same entropy bounds as the likelihood class used in Lemma 4. In particular, $\mathcal{J}_{n}$ is $\mathbb{P}$–Glivenko–Cantelli, so
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|\;\xrightarrow{P}\;0\,.$ (26)
Fix $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$ in $\mathcal{N}_{n}$. For each observation $\bf W$},
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|$ $=\Biggl\|\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\Bigl[\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr]\,dt\Biggr\|$
$\leq\int_{0}^{\tau}\|{\bf Z}{\bf Z}^{\top}\|\,Y(t)\Bigl|\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr|\,dt\,.$
Using the mean-value theorem for the exponential and boundedness of $\bf Z$} and $g$, there exists a constant $C>0$, depending only on model constants, such that
$\Bigl|\exp\{\mbox{${\theta}$}_{1}^{\top}{\bf Z}+g_{1}(t,{\bf X})\}-\exp\{\mbox{${\theta}$}_{2}^{\top}{\bf Z}+g_{2}(t,{\bf X})\}\Bigr|\;\leq\;C\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,.$
Therefore,
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|\;\leq\;C^{\prime}\int_{0}^{\tau}Y(t)\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,dt$
for some $C^{\prime}>0$. Taking expectations and using $\int_{0}^{\tau}Y(t)\,dt\leq\tau$,
$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|$ $\leq C^{\prime}\tau\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+C^{\prime}\,E_{0}\!\left\{\int_{0}^{\tau}Y(t)\,|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\,dt\right\}.$
By the Hölder regularity and boundedness assumptions on $g$, together with the fact that $d(\eta_{1},\eta_{2})$ controls the $L^{2}$–distance between $\chi_{\eta_{1}}$ and $\chi_{\eta_{2}}$, one can bound the right-hand side by $C^{\prime\prime}d(\eta_{1},\eta_{2})$ for some $C^{\prime\prime}>0$. In particular, for all $\eta\in\mathcal{N}_{n}$,
$\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\leq\;C^{\prime\prime}d(\eta,\eta_{o})\;\leq\;C^{\prime\prime}M\tau_{n}\;\longrightarrow\;0\,$,
since $\tau_{n}\to 0$. Hence
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\longrightarrow\;0\,.$ (27)
By Theorem 3, the efficient score for $\mbox{${\theta}$}_{o}$ is
$\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t)\,$,
and the efficient information matrix is
$I(\mbox{${\theta}$}_{o})=E_{0}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
By semiparametric information theory, the efficient score satisfies the usual information identity along the least favorable submodel. In particular, the efficient Fisher information equals minus the expectation of the derivative of the efficient score,
$\mathbb{P}\dot{s}(\mbox{${\theta}$}_{o},g_{o})=I(\mbox{${\theta}$}_{o})\,.$
Recalling that $\dot{s}(\mbox{${\theta}$}_{o},g_{o})=-J(\eta_{o};{\bf W})$, this gives
$\mathbb{P}J(\eta_{o};\cdot)=I(\mbox{${\theta}$}_{o})\,.$ (28)
Finally, from (26), (27), and (28),
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|$ $=\sup_{\eta\in\mathcal{N}_{n}}\bigl\|-\mathbb{P}_{n}J(\eta;\cdot)+I(\mbox{${\theta}$}_{o})\bigr\|$
@@
$\xrightarrow{P}0+0\;=\;0.$
This proves the lemma. ∎
@@
$\sup_{\eta\in\mathcal{N}_{n}}|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\;\leq\;C_{1}\delta_{n}\,$,
@@
$\sup_{\eta\in\mathcal{N}_{n}}\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\|\;\leq\;C_{2}\delta_{n}\,.$
@@
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl|\ell_{n}(\eta)-\widetilde{\ell}_{n}(\eta)-\{\ell_{n}(\eta_{o})-\widetilde{\ell}_{n}(\eta_{o})\}\bigr|=O_{p}(\delta_{n})\,$, (29)
@@
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\bigr\|=O_{p}(\delta_{n})\,.$ (30)
These bounds reflect that the per-subject approximation error is $O(\delta_{n})$ and the network outputs are uniformly bounded because the sieve $\mathcal{G}(K,s,\boldsymbol{p},D)$ is constructed with bounded weights.
By (29) and Theorem 1, on the ball $\{d(\eta,\eta_{o})\leq\tau_{n}\}$ we have
@@
$=O_{p}(\tau_{n}^{2})+O_{p}(\delta_{n})=O_{p}(\tau_{n}^{2}),$
[…]
@@
$S_{n}(\widetilde{\eta})-S_{n}(\widehat{\eta})=\underbrace{S_{n}(\widetilde{g},\widetilde{}\mbox{${\theta}$})-S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})}_{A}+\underbrace{S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})}_{B}.$
and (31) becomes
@@
By the mean value theorem in $\theta$}, there exists $\mbox{${\theta}$}^{\dagger}$ on the segment between $\widehat{}\mbox{${\theta}$}$ and $\widetilde{}\mbox{${\theta}$}$ such that
$A=\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\,(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\,.$
By Lemma 9, $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\xrightarrow{P}-I(\mbox{${\theta}$}_{o})$ and $\|\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}\|=O_{p}(1)$. Substituting this into (31), we obtain
@@
$\sqrt{n}\,R_{n}(\widetilde{\eta})=O_{p}(\sqrt{n}\delta_{n})=o_{p}(1),$
because $\sqrt{n}\delta_{n}\to 0$ by assumption.
@@
$B=S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})=\{\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\}+\{(\mathbb{P}_{n}-\mathbb{P})S(\widetilde{g},\widehat{}\mbox{${\theta}$})-(\mathbb{P}_{n}-\mathbb{P})S(\widehat{g},\widehat{}\mbox{${\theta}$})\},$
@@
$\|\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\|\lesssim\|\widetilde{g}-g_{o}\|_{L^{2}(\Delta)}^{2}+\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}^{2}=O_{p}(\tau_{n}^{2})\,.$
Moreover, Lemma 4 applied to the localized score class $\{S(g,\widehat{}\mbox{${\theta}$};\cdot):d((g,\widehat{}\mbox{${\theta}$}),\eta_{o})\leq C\tau_{n}\}$ implies
$\sup_{d((g,\widehat{\theta}),\eta_{o})\leq C\tau_{n}}|(\mathbb{P}_{n}-\mathbb{P})S(g,\widehat{}\mbox{${\theta}$};\cdot)|=O_{p}(\tau_{n}^{2})\,$,
and since both $(\widetilde{g},\widehat{}\mbox{${\theta}$})$ and $(\widehat{g},\widehat{}\mbox{${\theta}$})$ lie in this ball with probability 1o(1)1-o(1), we obtain
@@
because $\sqrt{n}\tau_{n}^{2}\to 0$. Using (32) together with $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}=O_{p}(1)$, $\sqrt{n}B=o_{p}(1)$, and $\sqrt{n}R_{n}(\widetilde{\eta})=o_{p}(1)$, we obtain

### v2

@@ Appendix A Proofs for Section 4
By Assumptions A2A3, the covariates are bounded and the sieve is uniformly bounded, so there exists a constant $C<\infty$ such that
$\sup_{\eta\in\mathcal{N}_{n}}\|J(\eta;{\bf W})\|\leq C\qquad\text{a.s.}$
Since $p$ is fixed, it suffices to control each matrix entry separately. The class
$\mathcal{J}_{n}=\{J(\eta;\cdot):\eta\in\mathcal{N}_{n}\}$
inherits the same bracketing entropy bound as the localized likelihood/score classes used in Lemma 10. Hence it is $\mathbb{P}$–Glivenko–Cantelli, and
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|\;\xrightarrow{P}\;0.$ (A.15)
Next we show continuity of $\eta\mapsto\mathbb{P}J(\eta;\cdot)$ in the metric $d$. Let $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$, and write $\phi(t,{\bf X},{\bf Z})=\chi_{\eta_{1}}(t,{\bf X},{\bf Z})-\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$. Then, by boundedness of $\bf Z$} and the mean-value theorem for the exponential, there exists a constant $C_{1}>0$ such that
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|$ $\leq C_{1}\int_{0}^{\tau}Y(t)\,|\phi(t,{\bf X},{\bf Z})|\,dt.$
Taking expectations and applying Cauchy–Schwarz,
$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|$ $\leq C_{1}\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)\,|\phi(t,{\bf X},{\bf Z})|\,dt\right]$
$\leq C_{1}\tau^{1/2}\left\{\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)\phi(t,{\bf X},{\bf Z})^{2}\,dt\right]\right\}^{1/2}.$
Under Assumptions A1A4, the true hazard $h_{o}(t\mid{\bf X},{\bf Z})=\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})\}$ is bounded away from zero, say $h_{o}\geq c_{h}>0$. Therefore,
$\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)\phi(t,{\bf X},{\bf Z})^{2}\,dt\right]$ $\leq c_{h}^{-1}\mathbb{E}_{0}\!\left[\int_{0}^{\tau}Y(t)h_{o}(t\mid{\bf X},{\bf Z})\phi(t,{\bf X},{\bf Z})^{2}\,dt\right]$
$=c_{h}^{-1}\mathbb{E}_{0}\!\left[\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\right]=c_{h}^{-1}d(\eta_{1},\eta_{2})^{2},$
where the equality follows from the compensator identity. Hence
$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|\leq C_{2}\,d(\eta_{1},\eta_{2})$
for some constant $C_{2}>0$. In particular,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\leq C_{2}M\tau_{n}\longrightarrow 0.$ (A.16)
At the true parameter,
$\mathbb{P}J(\eta_{o};\cdot)$ $=\mathbb{E}_{0}\!\left[\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})\}\,dt\right]$
$=\mathbb{E}_{0}\!\left[\Delta\,{\bf Z}{\bf Z}^{\top}\right]=I_{0}(\mbox{${\theta}$}_{o}),$
again by the compensator identity. Therefore, using (A.15) and (A.16),
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|$ $=\sup_{\eta\in\mathcal{N}_{n}}\bigl\|-\mathbb{P}_{n}J(\eta;\cdot)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|$
@@
$\xrightarrow{P}0.$
If $I_{0}(\mbox{${\theta}$}_{o})$ is nonsingular, let $\lambda_{0}=\lambda_{\min}(I_{0}(\mbox{${\theta}$}_{o}))>0$. On the event
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|<\lambda_{0}/2,$
all matrices $\dot{S}_{n}(\eta)$ are nonsingular and
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)^{-1}\bigr\|\leq 2/\lambda_{0}.$
Since this event has probability tending to one, we conclude
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)^{-1}\bigr\|=O_{p}(1).$
This completes the proof. ∎
@@
$\sup_{\eta\in\mathcal{N}_{n}}|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\;\leq\;C_{1}\omega_{n}(\delta_{n})\,$,
@@
$\sup_{\eta\in\mathcal{N}_{n}}\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\|\;\leq\;C_{2}\omega_{n}(\delta_{n})\,.$
@@
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl|\ell_{n}(\eta)-\widetilde{\ell}_{n}(\eta)-\{\ell_{n}(\eta_{o})-\widetilde{\ell}_{n}(\eta_{o})\}\bigr|=O_{p}\!\bigl(\omega_{n}(\delta_{n})\bigr)\,$, (A.17)
@@
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\bigr\|=O_{p}\!\bigl(\omega_{n}(\delta_{n})\bigr)\,.$ (A.18)
These bounds reflect that the per-subject approximation error is controlled by the local time-modulus $\omega_{n}(\delta_{n})$ from Lemma 14, while the network outputs are uniformly bounded because the sieve $\mathcal{G}(K,s,\boldsymbol{p},D)$ is constructed with bounded weights.
By (A.17) and Theorem 1, on the ball $\{d(\eta,\eta_{o})\leq\tau_{n}\}$ we have
@@
$=O_{p}(\tau_{n}^{2})+O_{p}\!\bigl(\omega_{n}(\delta_{n})\bigr)=O_{p}(\tau_{n}^{2}),$
because $\omega_{n}(\delta_{n})=o(\tau_{n}^{2})$ by assumption. The population loss satisfies $\ell_{0}(\eta)-\ell_{0}(\eta_{o})\asymp-d(\eta,\eta_{o})^{2}$ by Lemma 3. Therefore, the same M-estimation argument used in the proof of Theorem 1 for the oracle estimator (via Theorem 3.4.1 in van der Vaart and Wellner, 1996) applies to $\widetilde{\ell}_{n}$, yielding $d(\widetilde{\eta},\eta_{o})=O_{p}(\tau_{n})$ and $\|\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, which proves part 1.
@@
$S_{n}(\widetilde{\eta})-S_{n}(\widehat{\eta})=\underbrace{S_{n}(\widetilde{}\mbox{${\theta}$},\widetilde{g})-S_{n}(\widehat{}\mbox{${\theta}$},\widetilde{g})}_{A}+\underbrace{S_{n}(\widehat{}\mbox{${\theta}$},\widetilde{g})-S_{n}(\widehat{}\mbox{${\theta}$},\widehat{g})}_{B}\,.$
@@
$A=\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})\,(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\,.$
By Lemma 15, $\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})\xrightarrow{P}-I_{0}(\mbox{${\theta}$}_{o})$ and $\|\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})^{-1}\|=O_{p}(1)$. Substituting this into (A.19), we obtain
@@
$\sqrt{n}\,R_{n}(\widetilde{\eta})=O_{p}\!\bigl(\sqrt{n}\,\omega_{n}(\delta_{n})\bigr)=o_{p}(1),$
because $\sqrt{n}\,\omega_{n}(\delta_{n})\to 0$ by assumption.
@@
[…]
@@
$\|\mathbb{P}S(\widehat{}\mbox{${\theta}$},\widetilde{g})-\mathbb{P}S(\widehat{}\mbox{${\theta}$},\widehat{g})\|\lesssim\|\widetilde{g}-g_{o}\|_{L^{2}(\Delta)}^{2}+\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}^{2}=O_{p}(\tau_{n}^{2})\,.$
Moreover, Lemma 10 applied to the localized score class $\{S(\widehat{}\mbox{${\theta}$},g;\cdot):d((\widehat{}\mbox{${\theta}$},g),\eta_{o})\leq C\tau_{n}\}$ implies
$\sup_{d((\widehat{\theta},g),\eta_{o})\leq C\tau_{n}}|(\mathbb{P}_{n}-\mathbb{P})S(\widehat{}\mbox{${\theta}$},g;\cdot)|=O_{p}(\tau_{n}^{2})\,$,
[…]
@@
because $\sqrt{n}\tau_{n}^{2}\to 0$. Using (A.20) together with $\dot{S}_{n}(\mbox{${\theta}$}^{\dagger},\widetilde{g})^{-1}=O_{p}(1)$, $\sqrt{n}B=o_{p}(1)$, and $\sqrt{n}R_{n}(\widetilde{\eta})=o_{p}(1)$, we obtain